OpenStudy (anonymous):

A function is shown below: f(x) = x3 + 2x2 - x - 2 Part A: What are the factors of f(x)? Show your work. (3 points) Part B: What are the zeros of f(x)? Show your work. (2 points) Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). (5 points)

OpenStudy (solomonzelman):

Part A: Factor Part B: Finish solving after factoring Part C: Plug in any values for x, get the y values, re-write them as (x,y) coordinates and graph.

OpenStudy (anonymous):

are you doing y+mx=b

OpenStudy (anonymous):

Mm.. I don't think so.

OpenStudy (anonymous):

I think the factors are (x+1) (x-1) and (x+2) which would then mean the zeros are -1, 1, -2. And the end behavior would be that the graph falls to the left and rises to the right. But Im not sure how to describe how to graph,

jhonyy9 (jhonyy9):

so like a first step factoriz out and will get x^3 +2x^2 -x -2 = 0 x^2(x+2) -(x+2)=0 do xou can continue it ?

OpenStudy (anonymous):

No, I don't think so.

jhonyy9 (jhonyy9):

ok factoriz out the (x+2) and will get (x^2 -1)(x+2)=0

jhonyy9 (jhonyy9):

but you know again that x^2 -1 = (x-1)(x+1) yes ?

OpenStudy (anonymous):

So Im right?

jhonyy9 (jhonyy9):

yes the zeros are ok

OpenStudy (anonymous):

Yay! Okay, so then describing how to graph, does this sound okay? "To graph the f(x) you would plot your zeros, then the y-intercept, and draw a parabola through the points. In the end behavior, the graph falls to the left and rises to the right."

jhonyy9 (jhonyy9):

why say parabola ? when you know it sure that a function when has a graph of parabola ?when there is an quadratic equation yes ?

OpenStudy (anonymous):

Yes.

jhonyy9 (jhonyy9):

so than this mean that your graph will be marked from 3 lines ok ?

jhonyy9 (jhonyy9):

hope so much that you understand it now clearly

OpenStudy (anonymous):

Thanks, and I do! Your help is greatly appreciated :)

jhonyy9 (jhonyy9):

np was my pleasure good luck bye